A simple pendulum is found to vibrate 50 times within 200 s. When 1.5 m of its length is reduced to a certain length, it vibrates 50 times in 175 s. Find the original length of the pendulum.


1
Expert's answer
2021-12-02T10:08:39-0500

For original length:


"\frac{T_1}{N_1}=2\pi\sqrt{\frac{l}{g}},\\\space\\
l=g\bigg(\frac{T_1}{2\pi N_1}\bigg)^2=3.97\text{ s}."